OpenStudy (anonymous):

Solve for x. km/3-2x=4

OpenStudy (anonymous):

can i do this one

OpenStudy (anonymous):

wat is km exactly

OpenStudy (anonymous):

Is it kilometers

OpenStudy (anonymous):

x=km-4 ----- 2 x=4-km ----- 6 x=2-km ----- 3 x=km-12 ----- 6